OpenStudy (anonymous):

A bag has 4 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a yellow marble, do not replace it, and pick a red marble?

OpenStudy (anonymous):

2/15 1/15 1/9 1/10

hartnn (hartnn):

total how many marbles are there?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

is it d?

hartnn (hartnn):

probability of picking yellow(3) will be 3/10 since u are not replacing it ,now there are 9 marbles,to get red(3),the probability is 3/9 so,overall probab.=3/10*3/9=1/10 so yes d.

OpenStudy (anonymous):

thankss!

hartnn (hartnn):

welcome :)